OpenStudy (anonymous):

reduce the equation to quadratic form. 2x − 2/3 + x−1/3 + 6 = 0

OpenStudy (anonymous):

2/3+1/3 = 1 so it's 3x=-5 The rest you can do yourself.

OpenStudy (anonymous):

these are my options u = x−1/3 u = x1/3 u = x−2 u = x2 u = x−2/3 u = x3 u = x2/3 u = x4

OpenStudy (anonymous):

first one works, try it and see

OpenStudy (anonymous):

2x − 2/3 + x−1/3 + 6 = 0 is 3x+6=1 x+2-1/3=0 x=-(5/3) Where do you get u from?

OpenStudy (anonymous):

because i believe that the question is actually \[2x^{ − 2/3} + x^{−1/3} + 6 = 0\]otherwise it doesn't make any sense

OpenStudy (anonymous):

Oh